# A Sloppy approach to QSP: XAI enabling fit-for-purpose models

### v2

@@ Abstract
Quantitative Systems Pharmacology (QSP) promises to accelerate drug development, enable personalized medicine, and improve the predictability of clinical outcomes. Realizing its full potential depends on effectively managing the complexity of the underlying mathematical models and biological systems. Here, we present and validate a novel QSP workflow grounded in the principles of sloppy modeling, offering a practical and principled strategy for building and deploying models in a QSP pipeline. Our approach begins with a literature-derived model, constructed to be as comprehensive and unbiased as possible by drawing from the collective knowledge of prior research. At the core of the workflow is the Manifold Boundary Approximation Method (MBAM), which simplifies models while preserving their predictive capacity and mechanistic interpretability. Applying MBAM as a context-specific model reduction strategy, we link the simplified representation directly to the downstream predictions of interest. The resulting reduced models are computationally efficient and well-suited to key QSP tasks, including virtual population generation, experimental design, and target discovery. We demonstrate the utility of this workflow through case studies involving the coagulation cascade and SHIV infection. Our analysis suggests several promising next steps for improving the efficacy of bNAb therapies in HIV infected patients within the context of a general-purpose QSP modeling workflow.
@@
Quantitative Systems Pharmacology (QSP) uses mathematical models to understand complex biological systems and improve drug development. However, these models can become too complicated to work with effectively. We present a new modeling workflow that helps QSP scientists manage this complexity without sacrificing scientific rigor. Starting from models built entirely from published research, we use automated tools to simplify them based on specific questions, while preserving their ability to make reliable predictions. This approach supports key QSP tasks like virtual population generation, experimental design, and target selection. We show how it works on examples of blood coagulation and SHIV infection.
@@ 1 Introduction
Quantitative Systems Pharmacology (QSP) is an interdisciplinary field whose goals are to understand, predict, and control the behavior of complex biological systems[1, 2, 3]. Detailed mathematical and computational models are at the heart of the QSP approach that, distinct from traditional approaches to drug development, incorporates information about mechanisms to elucidate the biochemical foundation of drug efficacy, toxicity, and resistance at the system level. Such an approach incorporates diverse biological, pharmacological, and clinical data to create comprehensive models of drug action. These models promise to make predictions under novel conditions, accelerate and reduce risk in drug development, and inform patient-specific treatment decisions in personalized medicine[4, 5]. Because the biochemical processes underlying pharmacology are complex, many of the challenges faced by QSP modelers are direct consequences of the complexity inherent in the systems themselves, including computational cost, numerical stability and conditioning, data requirements, and unidentifiable parameters. Consequently, there is no consensus workflow for building and validating QSP models[6, 7, 8, 9], with many recent reviews emphasizing the need for techniques to address these issues within a QSP framework[10, 11, 12]. In this paper, we present a “sloppy model” workflow for QSP modeling built on rigorous mathematical and deep philosophical foundations based on decades of careful research that endows researchers with an understanding of key mechanisms in their otherwise impenetrably complex systems.
@@ 1.1 Existing Approaches to Managing Model Complexity
Effective strategies to address complexity in systems theory are informed by the modeling objectives. QSP models guide key decisions throughout the drug development process[4], often making predictions under novel conditions for which experimental data are not available and for which the relevance of diverse biological mechanisms is not known. Consequently, complexity in QSP models tends to accumulate, often manifesting as a proliferation of unidentifiable parameters and–plausible but indistinguishable –mechanistic alternatives. It is widely recognized that effectively navigating complex models involves trade-offs between physical realism and parsimony[13, 10, 7] with parameter identifiability being a central issue[14, 11]. Parameter identifiability analysis methods include those based on Fisher Information[15, 11] and profile likelihood[16, 15], while ensembles of unidentifiable parameters may be represented in virtual populations[17, 18, 19]. A powerful approach involves rewriting the model parameters in terms of the identifiable combinations[20]. Experimental design methods can determine maximally informative experiments within the limitations of the model[21, 22, 23, 24].
@@
Sensitivity analysis methods are also widely used in QSP to assess how model parameters influence system behavior. Local sensitivity analysis (LSA), such as those based on the Fisher Information Matrix, are computationally efficient but can be misleading when sensitivity varies significantly across parameter space. Global sensitivity analysis (GSA) techniques—such as Sobol’ indices [36, 37], Fourier-based methods (e/FAST)[38], and sampling-based approaches like Morris screening[39]—address this limitation by estimating parameter influence over a broad range of values, but they become computationally expensive in high-dimensional models.
@@
Importantly, both LSA and GSA can obscure the mechanistic origins of the sensitivities they quantify. A model may appear insensitive to certain parameters for reasons that range from biologically meaningful to mathematically trivial. For example, if parameters are expressed in poorly scaled or unnatural units, both LSA and GSA may over- or under-estimate their importance. In contrast, sloppy model analysis, described below, provides a coordinate-invariant framework for quantifying parameter importance based on the intrinsic geometry of the model manifold. This approach suggests a way forward that does more than suggest reductions: it certifies that the resulting reduced-order models retain predictive fidelity and reflect biologically meaningful simplifications, rather than artifacts of parameterization.
1.2 Sloppy Modeling as a Theoretical Foundation
The approach to managing complexity in QSP modeling presented here is based on techniques and ideas drawn from the so-called “Sloppy model” literature[40, 41, 42]. Sloppy modeling is a collection of techniques originating from statistical physics for analyzing multi-parameter models of complex systems. In this framework, parsimony is justified not because systems are inherently simple, but because most of the complexity is irrelevant for explaining or predicting a phenomenon of interest. A clear articulation of this philosophy has emerged from the statistical physics community over the last half century[43, 44, 45, 46, 47, 48, 49]. The sloppy model framework rigorously formalizes key concepts using information geometry, which further imbues it with powerful methodologies for model analysis, including those originating in biology[50, 51, 33]. Among the most powerful techniques of sloppy model analysis are model reduction methods for building fit-for-purpose parsimonious models that emphasize key mechanistic insights with mathematically rigorous performance guarantees[32, 41, 33]. The simplified models retain physical explainability of key mechanisms that control a phenomenon of interest, e.g. feedback loops. They have a more manageable computational cost compared to their complex progenitors. They are likewise numerically more stable and well-conditioned while clarifying the data requirements and parameters needed to make precise predictions for key quantities of interest. A sloppy modeling approach is therefore well suited for reasoning about relationships between mechanisms and systems behavior in QSP problems.
@@ 1.2 Sloppy Modeling as a Theoretical Foundation
In sloppy model analysis, a parameter’s relevance is how accurately that parameter must be known for the model to make a precise prediction as quantified by the Fisher Information Matrix (FIM). Parameter relevance is therefore context-dependent, i.e. conditional on the specific predictions the model is used to make. The term “sloppiness” was coined to describe a phenomenon first observed in systems biology models in which model parameters were largely unconstrained by available data, but those models nevertheless made precise predictions[40, 52]. In other words, models with unidentifiable parameters can make precise, falsifiable predictions, provided the unidentifiable parameters are also irrelevant to predictions of interest. For any predictions of interest, we can partition a model’s parameters into relevant and irrelevant combinations depending on whether they need to be finely tuned for the model to make precise predictions [^1] This binary partition of parameters as either relevant or irrelevant is a conceptual simplification. In practice, parameters exist on a continuum of relevance depending on how finely they must be tuned to make predictions at a target precision.. This perspective is bolstered by biological intuition in which phenomena of interest are often controlled by simple, understandable mechanisms of action embedded within the larger system, such as a key feedback loop or rate-limiting step. As long as the parameter combinations associated with relevant mechanisms are accurately tuned, the remaining parameter combinations can vary appreciably without impacting the model’s predictions. This observation suggests that the irrelevant parameters and the processes they parameterize can be ignored in the appropriate context while maintaining the relevant mechanistic features of the model.
@@
Of course, the challenges associated with complexity are not unique to QSP models; the search for parsimonious explanations is ubiquitous throughout science and engineering. Historically, scientific modeling has been the art of intuiting which mechanistic details are relevant and should therefore be included in the final model. Such insights are often developed by domain-specific experts whose entire careers are devoted to studying a few specific systems. Our most successful scientific theories back this intuition with mathematically rigorous abstractions. For example, continuum limits allow us to ignore positions of individual atoms or molecules and instead work with spatially averaged concentrations. In dynamical systems theory, singular perturbations abstract short-time dynamics and allow us to model only slower scales. Many other examples could be given. A major success of sloppy model analysis has been the systematic demonstration that in these historically well-understood cases, sloppy model analysis consistently reproduces conclusions derived by expert intuition from diverse scientific domains. However, it does so algorithmically by identifying the correct mathematical approximations that then guide intuition and new insights. Successes include phase transitions in statistical field theories[48], continuum limits (i.e., spatial averaging of large numbers of molecules)[48], singularly perturbed dynamical systems[53], balanced truncation in control systems[35, 34], and normal form theory in bifurcation analysis[54]. Accompanying these theoretical demonstrations are state-of-the-art technical tools for data fitting[50, 55], Bayesian sampling[56, 57], data visualization[58, 59], and parameter reduction of sloppy models[32, 33, 35]. Sloppy model analysis thus provides a suite of powerful computational tools built atop rigorous mathematical and philosophical foundations[41, 42]. These techniques collectively constitute a type of eXplainable Artificial Intelligence (XAI) that enable fit-for-purpose models in QSP and have the potential to support the QSP community in its goal of accelerating pharmaceutical development. By their nature, these tools should be applied judiciously depending on the specifics of the problem, and inform experts rather than replacing them.
1.3 Toward a Standardized Wokflow
@@
A primary objective of the present work is to formalize a QSP workflow based on the principles and tools of sloppy modeling. As QSP emerges as a tool for model-informed drug development, the need for standardized workflows becomes increasingly evident[60, 61]. The dual issues of model granularity and parameter estimation are universally recognized as a central, open problem[62, 63, 6, 64, 65]. In particular, balancing model complexity with parsimony can be challenging without a structured approach to scientific inquiry[63]. Several authors have proposed structured, iterative workflows to address this gap. Gadkar et al.[62] introduced a six-stage framework emphasizing goal definition, biological scoping, model development, and iterative evaluation. Helmlinger et al.[6] similarly highlighted the importance of cycles of refinement and validation grounded in empirical data. Tools like the QSP Toolbox[66] have been developed to operationalize these frameworks, supporting tasks such as model calibration and virtual population generation. Disease-specific applications, such as the workflow proposed by Go et al.[67] for atopic dermatitis trials, demonstrate how formalized QSP approaches can inform trial design and dosing strategies. A standardized QSP workflow enables model credibility, regulatory alignment, and cross-disciplinary communication.
@@
The workflow we propose is designed specifically to address issues related to model complexity in the broader QSP framework. As such, our proposed procedure complements most other workflows, and key elements can be incorporated into other operational paradigms. Key elements of our procedure are illustrated in Fig 1 and briefly summarized here. A more detailed discussion is given in section 2.1. Successful QSP flows from well-formulated questions; therefore, based on a specific QSP objective, e.g., target discovery, toxicology, etc., the modeling steps flow from this primary goal. First, a detailed, mechanistic QSP model is constructed by uncritically combining mechanisms found in the relevant literature, and its (many) parameters are estimated from available data. Next, specific modeling scenarios are designed, referred to as queries in our nomenclature (again, see section 2.1 for longer discussion). For example, one might be interested in predicting the system response under a novel drug regimen. These queries are typically scenarios for which data is unavailable but for which one would like to use the model to make predictions. Conceptually, the mathematical model is a machine that extracts information from the training data and propagates it forward to queries of interest. However, because of sloppiness, the relationship between information in fitting data and future predictions is obscured by the complexity of the literature-derived model, and many computational techniques are ill-conditioned. The next step, therefore, is to restrict the model to the parameters relevant to the queries. We construct a reduced order model based on these queries using a sloppy modeling algorithm called the Manifold Boundary Approximation Method (MBAM)[32]. By design, MBAM retains those parameters that are relevant to both the available data and the predictions of interest, clarifying the relationship between the two. The reduced order model determines what information was learned from the data and how it propagates forward to inform the QSP objectives. Critically, the model may not be identifiable from the data, but unidentifiable parameters in the reduced model reflect a fundamental limitation in the information available in the data. Finally, subsequent QSP calculations–such as hypothesis testing, virtual population generation, or experimental design–are performed using the reduced model, and are informed by which modeling elements were retained by the reduction process. In our experience, nearly all practical modeling tasks are improved, accelerated, and clarified by working with the parsimonious reduced model.
@@
Fig 1: A Sloppy Workflow for QSP. After building detailed, literature-derived models, parameter values are calibrated by fitting to system data. Carefully selecting “model queries” based on the target application, a reduced-order model is built that removes sloppy parameters and models the “active manifold” that efficiently propagates information from available data to predictions of interest. Calculations performed on the reduced model are mechanistically interpretable while having improved numerical conditioning, stability, and computational cost.

### v3

# A CRISP approach to QSP: XAI enabling fit-for-purpose models

@@ Abstract
Quantitative Systems Pharmacology (QSP) promises to accelerate drug development, enable personalized medicine, and improve the predictability of clinical outcomes. Realizing this potential requires effectively managing the complexity of mathematical models representing biological systems. Here, we present and validate a novel QSP workflow–CRISP (Contextualized Reduction for Identifiability and Scientific Precision)–that addresses a central challenge in QSP: the problem of complexity and over-parameterization, in which models contain irrelevant parameters that obscure interpretation and hinder predictive reliability. The CRISP workflow begins with a literature-derived model, constructed to be comprehensive and unbiased by integrating prior mechanistic insights. At the core of the workflow is the Manifold Boundary Approximation Method (MBAM), a reduction technique that simplifies models while preserving mechanistic structure and predictive fidelity. By applying MBAM in a context-specific manner, CRISP links parsimonious models directly to predictions of interest, clarifying causal structure and enhancing interpretability. The resulting models are computationally efficient and well-suited to key QSP tasks, including virtual population generation, experimental design, toxicology, and target discovery. We demonstrate the utility of CRISP on case studies involving the coagulation cascade and SHIV infection, and identify promising directions for improving the efficacy of bNAb therapies for HIV. Together, these results establish CRISP as a general-purpose QSP workflow for turning complex mechanistic models into tools for precise scientific reasoning to guide pharmacological and regulatory decision-making.
@@
Quantitative Systems Pharmacology (QSP) uses mathematical models to understand complex biological systems and support drug development. However, these models are often so complex that they limit precise scientific reasoning. We introduce a new modeling workflow—CRISP (Contextualized Reduction for Identifiability and Scientific Precision)—designed to help scientists manage this complexity while improving the precision and reliability of QSP analyses. Starting from comprehensive models built from published research, CRISP uses automated tools to simplify models based on specific scientific questions, preserving their ability to make accurate predictions. This pipeline supports core QSP applications, including virtual population generation, experimental design, and target selection. We demonstrate CRISP using case studies in blood coagulation and SHIV infection.
@@ 1 Introduction
Quantitative Systems Pharmacology (QSP) is an interdisciplinary field whose goals are to understand, predict, and control the behavior of complex biological systems[1, 2, 3]. Detailed mathematical and computational models are at the heart of the QSP approach that, distinct from traditional approaches to drug development, incorporates information about mechanisms to elucidate the biochemical foundation of drug efficacy, toxicity, and resistance at the system level. Such an approach incorporates diverse biological, pharmacological, and clinical data to create comprehensive models of drug action. These models promise to make predictions under novel conditions, accelerate and reduce risk in drug development, and inform patient-specific treatment decisions in personalized medicine[4, 5]. Because the biochemical processes underlying pharmacology are complex, many of the challenges faced by QSP modelers are direct consequences of the complexity inherent in the systems themselves, including computational cost, numerical stability and conditioning, data requirements, and unidentifiable parameters. Consequently, there is no consensus workflow for building and validating QSP models[6, 7, 8, 9], with many recent reviews emphasizing the need for techniques to address these issues within a QSP framework[10, 11, 12]. In this paper, we present and validate a novel QSP workflow–CRISP (Contextualized Reduction for Identifiability and Scientific Precision)–that addresses a central challenge in QSP: the problem of complexity and over-parameterization, in which models contain irrelevant parameters that obscure interpretation and hinder predictive reliability. The pipeline is built on rigorous mathematical and deep philosophical foundations based on decades of careful research into the theory of predictive modeling using Information Geometry that endows researchers with an understanding of key mechanisms in their otherwise impenetrably complex systems.
@@ 1.1 Existing Approaches to Managing Model Complexity
Effective strategies to address complexity in systems theory are informed by the modeling objectives. QSP models guide key decisions throughout the drug development process[4], often making predictions under novel conditions for which experimental data are not available and for which the relevance of diverse biological mechanisms is not known. Consequently, complexity in QSP models tends to accumulate, often manifesting as a proliferation of unidentifiable parameters and plausible – but indistinguishable –mechanistic alternatives. It is widely recognized that effectively navigating complex models involves trade-offs between physical realism and parsimony[13, 10, 7] with parameter identifiability being a central issue[14, 11]. Parameter identifiability analysis methods include those based on Fisher Information[15, 11] and profile likelihood[16, 15], while ensembles of unidentifiable parameters may be represented in virtual populations[17, 18, 19]. A powerful approach involves rewriting the model parameters in terms of the identifiable combinations[20]. Experimental design methods can determine maximally informative experiments within the limitations of the model[21, 22, 23, 24].
@@
Sensitivity analysis methods are also widely used in QSP to assess how model parameters influence system behavior. Local sensitivity analysis (LSA), such as those based on the Fisher Information Matrix, are computationally efficient but can be misleading when sensitivity varies significantly across parameter space. Global sensitivity analysis (GSA) techniques—such as Sobol’ indices [36, 37], Fourier-based methods (e/FAST)[38], and sampling-based approaches like Morris screening[39]—address this limitation by estimating parameter influence over a broad range of values, but they become prohibitively computationally expensive in high-dimensional models.
@@
Importantly, both LSA and GSA can obscure the mechanistic origins of the sensitivities they quantify. A model may appear insensitive to certain parameters for reasons that range from biologically meaningful to mathematically trivial. For example, if parameters are expressed in poorly scaled or unnatural units, both LSA and GSA may over- or under-estimate their importance. In contrast, Information Geometry analysis, described below, provides a coordinate-invariant framework for quantifying parameter importance based on the intrinsic geometry of the model manifold. This approach suggests a way forward that does more than suggest reductions: it certifies that the resulting reduced-order models retain predictive fidelity and reflect biologically meaningful simplifications, rather than artifacts of parameterization.
1.2 Information Geometry: A Theoretical Foundation for QSP
The approach to managing complexity in QSP modeling presented here draws on ideas and techniques from the so-called sloppy model literature [40, 41, 42]. The term sloppiness was first used to describe a phenomenon observed in systems biology models, where most parameters were poorly constrained by available data, yet the models nevertheless yielded reliable, precise predictions [40, 43]. This counter intuitive behavior arises when predictions depend primarily on a few combinations of parameters, rendering most parameters effectively irrelevant to the outputs of interest. Today, the term “sloppiness” remains as a misnomer, since mathematically rigorous methodologies based on Information Geometry now enable precise, mechanistic reasoning about the predictive capabilities of such models. We refer to these collection of theories and techniques as “Information Geometric analysis” to emphasize the rigorous foundation on which they are built.
In our Information Geometry framework, parsimony is not a reflection of inherent simplicity in biological systems, but rather a pragmatic response to the fact that much of a model’s complexity is irrelevant for answering specific scientific questions. This perspective has deep roots in statistical physics, where a rich tradition of theoretical work justifies simplified descriptions of complex systems [44, 45, 46, 47, 48, 49, 50]. Information geometry provides a rigorous foundation for identifying which aspects of a model matter most. This foundation supports a suite of powerful techniques, including biologically grounded model reduction methods that yield parsimonious, fit-for-purpose models with rigorous performance guarantees [51, 52, 32, 41, 33]. These reduced models retain key mechanistic interpretability—such as the role of feedback loops—while offering lower computational cost, greater numerical stability, and clearer parameter identifiability and data requirements. As such, Information Geometry provides an ideal foundation for precise reasoning about the links between biological mechanisms and system-level behavior in QSP applications.
@@ 1.2 Information Geometry: A Theoretical Foundation for QSP
In an Information Geometry analysis, a parameter’s relevance is how accurately that parameter must be known for the model to make a precise prediction, as quantified by the Fisher Information Matrix (FIM). Critically, parameter relevance is context-dependent, i.e. conditional on the specific predictions the model is used to make. For any predictions of interest, we can partition a model’s parameters into relevant and irrelevant combinations depending on whether they need to be finely tuned for the model to make precise predictions [^1] This binary partition of parameters as either relevant or irrelevant is a conceptual simplification. In practice, parameters exist on a continuum of relevance depending on how finely they must be tuned to make predictions at a target precision.. This perspective is bolstered by biological intuition: phenomena of interest are often controlled by simple, understandable mechanisms of action embedded within the larger system, such as a key feedback loop or rate-limiting step. As long as the parameter combinations associated with relevant mechanisms are accurately tuned, the remaining parameter combinations can vary appreciably without impacting the model’s predictions. This observation suggests that the irrelevant parameters and the processes they parameterize can be ignored in the appropriate context while maintaining the relevant mechanistic features of the model.
@@
Of course, the challenges associated with complexity are not unique to QSP models; the search for parsimonious explanations is ubiquitous throughout science and engineering. Historically, scientific modeling has been the art of separating relevant mechanistic details from irrelevant processes to arrive at the final model. Such insights are often developed by domain-specific experts whose entire careers are devoted to studying a few specific systems. Our most successful scientific theories back this intuition with mathematically rigorous abstractions. For example, continuum limits allow us to ignore positions of individual atoms or molecules and instead work with spatially averaged concentrations. In dynamical systems theory, singular perturbations abstract short-time dynamics and allow us to accurately model slower processes. Many other examples could be given. A major success of Information Geometry analysis has been the systematic demonstration that in these historically well-understood cases, Information Geometry consistently reproduces conclusions derived by expert intuition from diverse scientific domains. However, it does so algorithmically by identifying the correct mathematical approximations that then guide intuition and new insights. Successes include phase transitions in statistical field theories[49], continuum limits (i.e., spatial averaging of large numbers of molecules)[49], singularly perturbed dynamical systems[53], balanced truncation in control systems[35, 34], and normal form theory in bifurcation analysis[54]. Accompanying these theoretical demonstrations are state-of-the-art technical tools for data fitting[51, 55], Bayesian sampling[56, 57], data visualization[58, 59], and parameter reduction[32, 33, 35]. Information Geometry thus provides a suite of powerful computational tools built atop rigorous mathematical and philosophical foundations[41, 42]. These techniques collectively constitute a type of eXplainable Artificial Intelligence (XAI) that enable fit-for-purpose models in QSP and have the potential to support the QSP community in its goal of accelerating pharmaceutical development. By their nature, these tools should be applied judiciously depending on the specifics of the problem, and inform experts rather than replacing them.
1.3 Toward a CRISP Workflow
@@
A primary objective of the present work is to formalize a QSP workflow based on the principles and tools of Information Geometry. As QSP emerges as a tool for model-informed drug development, the need for standardized workflows becomes increasingly evident[60, 61]. The dual issues of model granularity and parameter estimation are universally recognized as a central, open problem[62, 63, 6, 64, 65]. In particular, balancing model complexity with parsimony can be challenging without a structured approach to scientific inquiry[63]. Several authors have proposed structured, iterative workflows to address this gap. Gadkar et al.[62] introduced a six-stage framework emphasizing goal definition, biological scoping, model development, and iterative evaluation. Helmlinger et al.[6] similarly highlighted the importance of cycles of refinement and validation grounded in empirical data. Tools like the QSP Toolbox[66] have been developed to operationalize these frameworks, supporting tasks such as model calibration and virtual population generation. Disease-specific applications, such as the workflow proposed by Go et al.[67] for atopic dermatitis trials, demonstrate how formalized QSP approaches can inform trial design and dosing strategies. A standardized QSP workflow enables model credibility, regulatory alignment, and cross-disciplinary communication.
@@
We introduce a workflow leveraging Contextualized Reduction for Identifiability and Scientific Precision: CRISP. The CRISP workflow is designed specifically to address issues related to model complexity and over-parameterization in the broader QSP framework. As such, CRISP complements most other workflows, and key elements can be incorporated into other operational paradigms. Key elements of CRISP procedure are illustrated in Fig 1 and briefly summarized here. A more detailed discussion is given in section 2.1. Successful QSP flows from well-formulated questions; therefore, based on a specific QSP objective, e.g., target discovery, toxicology, etc., the modeling steps flow from this primary goal. First, a detailed, mechanistic QSP model is constructed by uncritically combining mechanisms found in the relevant literature, and its (many) parameters are estimated from available data. Next, specific modeling scenarios are designed, referred to as queries in our nomenclature (again, see section 2.1 for longer discussion). For example, one might be interested in predicting the system response under a novel drug regimen. These queries are typically scenarios for which data is unavailable but for which one would like to use the model to make predictions. Conceptually, the mathematical model is a machine that extracts information from the training data and propagates it forward to queries of interest. However, the relationship between information in fitting data and future predictions is often obscured by the complexity of the literature-derived model, rendering many computational techniques ill-conditioned. The next step, therefore, is to restrict the model to the parameters relevant to the queries. We construct a reduced order model based on these queries using a Information Geometry algorithm called the Manifold Boundary Approximation Method (MBAM)[32]. By design, MBAM retains those parameters that are relevant to both the available data and the predictions of interest, clarifying the relationship between the two. The reduced order model determines what information was learned from the data and how it propagates forward to inform the QSP objectives. Critically, the model may not be identifiable from the data, but unidentifiable parameters in the reduced model reflect a fundamental limitation in the information available in the data. Finally, subsequent QSP calculations–such as hypothesis testing, virtual population generation, or experimental design–are performed using the reduced model, and are informed by which modeling elements were retained by the reduction process. In our experience, nearly all practical modeling tasks are improved, accelerated, and clarified by working with the parsimonious reduced model.
@@
Fig 1: Contextualized Reduction for Identifiability and Scientific Precision (CRISP) for QSP. After building a detailed, literature-derived model, parameter values are calibrated by fitting to system data. Carefully selecting “model queries” based on the target application, a reduced-order model is built that removes irrelevant parameters and models the “active manifold” that efficiently propagates information from available data to predictions of interest. Calculations performed on the reduced model are mechanistically interpretable while having improved numerical conditioning, stability, and computational cost.